OpenStudy (anonymous):

Kim is studying the sale of a particular brand of cereals from the year 2000 to 2010. She writes the following function to model the sale of the cereal S(t), in million dollars, after t years: S(t) = t2 + 7t + 69 Part A: What does the y-intercept of the graph of the function represent? (4 points) Part B: What is the reasonable domain of the graph of the function? (3 points) Part C: What is the average rate of change of the sale of the cereal from the first year to the fourth year? Show your work. (3 points)

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@JonnyVonny

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

@Stormswan

OpenStudy (anonymous):

I just need a reasonable domain

OpenStudy (anonymous):

and the average rate of change

OpenStudy (anonymous):

the y intercept I know

ganeshie8 (ganeshie8):

how many years are there between 2000 and 2010 ?

OpenStudy (anonymous):

10

ganeshie8 (ganeshie8):

0->10 thats 11 years, but the number as such doesn't matter much

OpenStudy (anonymous):

So 0=>10?

ganeshie8 (ganeshie8):

can we say the reasonable domain is from t=0 to t=10 ?

ganeshie8 (ganeshie8):

yes you have it !

OpenStudy (anonymous):

Ok thank you are you sure? and one more What is the average rate of change of the sale of the cereal from the first year to the fourth year? Show your work.

ganeshie8 (ganeshie8):

acerage rate of change is same as `slope`

ganeshie8 (ganeshie8):

start by evaluating the function at t=1 and t=4

ganeshie8 (ganeshie8):

S(t) = t^2 + 7t + 69 S(1) = ? S(4) = ?

OpenStudy (anonymous):

So 4-1=3?

ganeshie8 (ganeshie8):

not quite

ganeshie8 (ganeshie8):

you need to evaluate the function at those t values and plug them in below formula : \[\large\text{Average rate of change} = \dfrac{S(4) - S(1)}{4-1}\]

OpenStudy (anonymous):

1=77 4=113

OpenStudy (anonymous):

12?

ganeshie8 (ganeshie8):

looks good!

OpenStudy (anonymous):

Ok are you sure about the domain part?

ganeshie8 (ganeshie8):

aren't you sure ?

OpenStudy (anonymous):

No I am just asking, you the math proffessor/expert

OpenStudy (anonymous):

becuase when I graph it it doesn't quite make sense

OpenStudy (anonymous):

@ganeshie8 ?

ganeshie8 (ganeshie8):

could you elaborate on what specific thing that doesn't make sense ?

OpenStudy (anonymous):

Ok, when You graph S(t) = t^2 + 7t + 69 you get a parabola, an the domain just doesn't make sense in the graph

ganeshie8 (ganeshie8):

domain is x values right ?

ganeshie8 (ganeshie8):

look at the graph between t=0 and t=10

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

the graph starts at t=0, this represents the year 2000, and the value is 69 million dollars

OpenStudy (anonymous):

ok, and how does it end on 10?

ganeshie8 (ganeshie8):